You have a 750mL of a buffer solution that is composed of NH3 and NH4Cl with a Kb value 1.8x 10^5 and a pH 9.43...... How many mL 6.00M HCl can be added before the buffer is exhausted?
Answer #1

this is a NH3-NH4+ buffer

pKb = -log(1.8*10^-5) = 4.75

pKa = 14-4.75= 9.25

then..

pH = pKa + log(NH3/NH4+)

9.43 = 9.45 + log(NH3/NH4+)

10^(9.43 - 9.45 ) = (NH3/NH4+)

0.95499 = (NH3/NH4+)

assume total NH3 + NH4+ mix is = 1 M then

mol of NH3 + NH4+ = 0.75 mol

0.95499 = (NH3/NH4+)

NH3 = 0.95499 *NH4+

0.95499 *NH4++ NH4+ = 0.75

NH4+ = 0.75/(0.95499 +1) = 0.3836 mol

mol of NH3 --> 0.95499 *0.3836 = 0.36633

then

max amount of HCl possible --> mol of NH3

0.36633 mol of HCl

V = mol/M = (0.36633)/(6) = 0.061 = 61 mL max...

this assume total molarity of buffer = 1 M